OpenStudy (anonymous):

identify the conic and write in standard form x^2-9y^2+18y=36

OpenStudy (anonymous):

Hyperbola. x^2 - 9y^2 + 18y = 36 x^2 - 9(y^2 - 2y + 1) = 27 x^2 27 - (y - 1)^2/3 = 1